OpenStudy (anonymous):

Solve the equation x-8=16 and state which axiom do you use here

OpenStudy (anonymous):

x-8=16 add 8 to both sides: x-8+8=16+8 simplify x = 24 Have no idea what "axiom" that might be.

OpenStudy (anonymous):

Associative property maybe?

OpenStudy (anonymous):

Gawd...this gets goofy sometimes. I googled "what do you call adding a number to both sides" and damned if I didn't get some hits. Apparently this falls under "properties of equality", and specifically subtitled "subtraction property of equality" and the "addition property of equality" I don't know if that's what you're looking for or not, but there it is. :-)